OpenStudy (anonymous):

find laplace of f(t) = 3sin(5t + 45) for t >= 0

OpenStudy (anonymous):

Using the definition is probably your best bet.

OpenStudy (anonymous):

\[\mathcal{L}\left\{3\sin(5t+45)\right\}=3\int_0^\infty \sin(5t+45)~e^{-st}~dt\] Integrate by parts from here.

OpenStudy (anonymous):

are you sure coz i get confused when to use step function (heaviside function)

OpenStudy (anonymous):

coz in this previous problem "f sub 1(t) = 0 for t < 0" i used step function

OpenStudy (anonymous):

@Preetha @e.mccormick @AravindG